OpenStudy (anonymous):

1. You work at the Peawingspan Blue store 40 hours/week over five days at a rate of $9.25/hour. Deductions are FICA (7.65%), federal withholding (12%), and state withholding (8%). Your expenses are transportation at $5.25/day, lunch at $3.85/day, and black slacks required for work at $19.95. How much is your discretionary income for one work week? A. $324.50 B. $304.55 C. $222.19 D. $202.24

OpenStudy (anonymous):

Tiny bit more intense than the last one. Haha

OpenStudy (anonymous):

Right so first all you have to do is take the important information out.

OpenStudy (anonymous):

Alright.

OpenStudy (anonymous):

40 hours/week over five days at a rate of $9.25/hour. Deductions are FICA (7.65%), federal withholding (12%) state withholding (8%). transportation at $5.25/day, lunch at $3.85/day black slacks required for work at $19.95.

OpenStudy (anonymous):

Right, now we find the total income, so hours x rate then add the percentages of the taxes and stuff and add the per day payments for necessities

OpenStudy (anonymous):

and for now we will leave the slacks

OpenStudy (anonymous):

Okay, a moment.

OpenStudy (anonymous):

Take your time

OpenStudy (anonymous):

xD I screwed up somewhere. Not really sure how to add the percentages of the taxes an stuff.

OpenStudy (anonymous):

8+12+7.65=Total%

OpenStudy (anonymous):

2.61%?

OpenStudy (anonymous):

Hmm, just add the numbers i just gave you and put a percentage sign at the end of it. Out of curiosity how did you get 2.61

OpenStudy (anonymous):

or 9.56%?

OpenStudy (anonymous):

Oh shoot, that looked like a division sign. My bad.

OpenStudy (anonymous):

Oh i see :)

OpenStudy (anonymous):

27.65%

OpenStudy (anonymous):

xD haha

OpenStudy (anonymous):

Correct Total Income=? Total Percentage=27.65% And per day expenses=? 2 to go

OpenStudy (anonymous):

I'm to tired to figure Total income correctly, I tried multiple ways and it seems to be wrong every time. But per day expenses would be: 45.5

OpenStudy (anonymous):

Okayy, for income just do: Hours of work x Rate per hour. while i check the per day expenses

OpenStudy (anonymous):

370

OpenStudy (anonymous):

See, I was like dividing n stuff.

OpenStudy (anonymous):

lol, So per day expenses is correct, but you did it for the 5 days he works which is good to see so T.I. : $270 Tax:27.65 Expenses for 5 days of work: $45.5 Slacks: $19.95 Now, find 27.65% of 270

OpenStudy (anonymous):

What is the formula/"equation format" for that? And how did you get 270 for total income? Was I not supposed to do it for the 5 days he works?

OpenStudy (anonymous):

It said he works 40 hours per week, on which he works five days, 40 hours is already done according to 5 days

OpenStudy (anonymous):

27.65% of 270=74.655 Working example: Find 60% of 1200 To work out a percentage of something divide the percentage by 100 60 / 100 = 0.6 then multiply the answer by the subject number 0.6 * 1200 = 720 Answer: 60% of 1200 = 720

OpenStudy (anonymous):

Oooh. Okay.

OpenStudy (anonymous):

Now do we just deduct the expenses and slacks from the new total income?

OpenStudy (anonymous):

Right so now we add the all the costs together 74.655+45.5+19.95=?

OpenStudy (anonymous):

140.105

OpenStudy (anonymous):

370-140.105=229.9 Not one of the answers lemme try one more thing. Umm your addition of total costs is incorrect

OpenStudy (anonymous):

my bad it is right

OpenStudy (anonymous):

Maybe we can retrace our steps and plug in one of the multiple choice answers such as, $222.19, somehow and figure it out that way?

OpenStudy (anonymous):

okayy

OpenStudy (anonymous):

there are two things which i see where we might have gone wrong, either we have to take each percentage out separately or the slacks aren't a necessity

OpenStudy (anonymous):

Well it says the slacks are required for work, so I think they might be. Might have to take out each percent separately.

OpenStudy (anonymous):

Like you were saying.

OpenStudy (anonymous):

kayy lets do that then

OpenStudy (anonymous):

Oh oh, I think I know

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

Nvm, checked, not the issue.

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Be right back a second.

OpenStudy (anonymous):

Okay back

OpenStudy (anonymous):

7.65% of 370 = 28.305 370-28.305= 341.695 12% of 341.695= 41.0034 341.695 - 41.0034= 300.6916 8% of 300.6916 = 24.055328 300.6916 - 24.055328= 276.636272 276.636272-(45.5+19.95) =211.186272

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

So most likely 222.19?

OpenStudy (anonymous):

Id say so its in the middle of both answers

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

Thank you very much with this one especially, haha

OpenStudy (anonymous):

Well that took longer than i thought it would :P , sorry bout my long methods but i try to tell you how to do it so next time you dont have to come back here :)

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

Haha yeah, it is nice to learn it at times and not be told all the time. Although being told is perfectly fine, but I think I prefer this method, when I have more patience. It was D sadly.